OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x2 − 6x − 36.

OpenStudy (anonymous):

"using complete sentences" lol

OpenStudy (anonymous):

\[2x^2-6x-36\]First notice that there is a common factor of two.\[2(x^2-3x-18)\]Now we are looking for 2 numbers that add up to -3 and multiply to make -18. Namely -6 and +3\[2((x-6)(x+3\]

OpenStudy (anonymous):

I don't know why my LaTeX isn't working but, does that make sense?

OpenStudy (anonymous):

\[2((x-6)(x+3))\]

OpenStudy (unklerhaukus):

the other method is to use the quadratic formula \[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[(x-x_1)(x-x_2)=0\]

OpenStudy (anonymous):

Ok for factoring ac=-72 and b=6 so -12 and 6 are the numbers ac is the product and b is the sum of two numbers such that they satisfy both conditions. 2x^2-12x+6x-36=0 2x(x-6)+6(x-6)=0 (2x+6)(x-6)=0 thus x=-3 or x=6 Hence solved